OpenStudy (jenniferjuice):

Factor completely 5c5 + 60c4 + 180c3. 5c3(c + 6)(c - 6) 5c3(c2 + 12c + 36) 5c3(c + 6)2 5(c + 6)2

OpenStudy (solomonzelman):

\(\Large\color{darkgreen}{ 5c^5 + 60c^4 + 180c^3 }\) \(\Large\color{darkgreen}{ (5c^3)c^2 + (5c^3)12c + (5c^3)36 }\) \(\Large\color{darkgreen}{ 5c^3(c^2 + 12c +36 ) }\)

OpenStudy (solomonzelman):

Now, you need to factor the \(\Large\color{darkgreen}{ c^2 + 12c +36 }\)

OpenStudy (jenniferjuice):

5c3(c + 6)(c - 6)?

OpenStudy (solomonzelman):

no.... HINT: `c²+12c+36` is a perfect square.

OpenStudy (jenniferjuice):

wait so its just 5c3(c2 + 12c + 36) ?

OpenStudy (solomonzelman):

No it is not

OpenStudy (precal):

actually you can still factor the second part

OpenStudy (precal):

Think what 2 numbers multiplied give you 36 but when you add together give you 12

OpenStudy (jenniferjuice):

6 and 6

OpenStudy (precal):

yes but what are your signs + + - - + - -+

OpenStudy (jenniferjuice):

+6 and +6

OpenStudy (precal):

yes now which is it? recall c times c gives you c^2 6 times 6 gives you 36 +6+6 gives you the middle term of +12 (c+6)(c+6) gives you c^2+12c+36

OpenStudy (jenniferjuice):

5c3(c + 6)2?

OpenStudy (precal):

yes good job

OpenStudy (anonymous):

However, you would not cube the 5c in front of your (c+6)

OpenStudy (jenniferjuice):

so 5(c + 6)2?

OpenStudy (anonymous):

Oh sorry! my fault I made a mistake, you were right. I apologize

OpenStudy (anonymous):

\[5c^3(c+6)^2\] is correct

OpenStudy (anonymous):

i apologize

OpenStudy (precal):

the last term is the number you need so in this case it is 36 you list all of the factors of 36 1 times 36 2 times 18 3 times 12 4 times 9 6 times 6 the sign in front of the last term tells you to add all of the terms 1+36=37 2+18=20 3+12=15 4+9=13 6+6=12 sign also, tells you that your sign choices are - - or + + -6-6=-12 +6+6=+12 we need a +12 for the middle term put it all together (c+6)(c+6)